OpenStudy (anonymous):

Find the volume of the solid generated by revolving the described region about the given axis: The region bounded above by the line y=4 , below by the curve y=x^2 , rotated along the following lines: y=4

OpenStudy (anonymous):

\[\pi \int\limits_{-2}^{2}(x^2 - 4 ) ^2 dx\] If im not mistaken :) ( Im a bit rusty )

OpenStudy (anonymous):

Acutally π∫(4- x^2)2dx because y=4 > x^2 in the bounded area